OpenStudy (anonymous):

factor completly 81x4 -16

OpenStudy (anonymous):

Did you mean \(81x^4-16\)?

OpenStudy (anonymous):

Yes you are correct. I wrote it wrong.

OpenStudy (anonymous):

81 x 4 = 324 -16 = 308

OpenStudy (anonymous):

wait nevermind >;p

OpenStudy (anonymous):

Not wrong, you just didn't use the equation editor. It looks like the difference of two perfect squares to me.

OpenStudy (anonymous):

@gypsy1274 can you help me with 2 math questions ?

OpenStudy (anonymous):

I can try. Post them in your own thread, tag me and I'll look.

OpenStudy (anonymous):

\( (9x^2 - 4)(9x^2+4)=(3x-2)(3x+2)(9x^2+4)\)